Let (Z1, Y1), . . . , (Zn, Yn) be generated as follows:

Zi ∼ Bernoulli(p)

Yi ∼ { N(0, 1) if Zi = 0 , N(5, 1) if Zi = 1

(a) Assume we do not observe the Zi ’s. Write the pdf f(y) of Y as a mixture of two normal distribution pdf. (Use the notation φ(·) which is the standard normal pdf.)

(b) Write down the likelihood function for p (without Zi ’s).

(c) Write down the complete likelihood function for p (assuming the Zi ’s are observed).

(d) Find the maximum likelihood estimation of p using the likelihood from (c). 

1
Expert's answer
2021-05-13T08:40:51-0400

we see that N(0,1) and N(5,1) with probability 1-p and p respectively.


Thus density is "R(y)=(1-p)\dfrac{1}{\sqrt{2\pi}}e^{-\frac{y^2}{2}}+p\dfrac{1}{\sqrt{2\pi}}e^{-\frac{(y-5)^2}{2}}"


(b) The likelihood function from the above density, the likelihood function without the z's is:


"L_n(P)= ((1-p)\dfrac{1}{\sqrt{2\pi}}e^{-\frac{y_i^2}{2}}+p\dfrac{1}{\sqrt{2\pi}}e^{-\frac{(y_i-5)^2}{2}}"


(c) Note that if the "z_i" is known, then the likelihood for p depend exclusively on "Z_i" as y/z has a normal distribution independent of p.


Thus the complete likelihood function is-


"L_n(P)= ((1-p)\dfrac{1}{\sqrt{2\pi}}e^{-\frac{y_i^2}{2}}+p\dfrac{1}{\sqrt{2\pi}}e^{-\frac{(y_i-5)^2}{2}}"


.


(d) As we know, The likelihood for p depends exclusively on "Z_i's \text{ as } Y|Z" has a normal distribution of p.


Thus the likelihood of p now is

 "L_n(p)=f_n(Y,Z|p)=f_n(Z|p)\times f_n(Y|Z)," where "f_n" is the joint density of the sample. Here "f_n(Y|Z)" is independent of p and thus


     "L_n(p) \propto f_n(Z|p)"


But now, note that "Z_1,Z_2,...,Z_n"


"\Rightarrow f_n(Z|p)=\Pi_{i=1}^n P^{Z_i}(1-p)^{1-Z_i}"

'

         "=p^{\sum Z_i}(1-p)^{n-\sum Z_i}"


Thus the likelihood function is-


"L_n(p) \propto p^{\sum Z_i}(1-p)^{n-\sum Z_i}"


Taking log of above equation and we get-


"LogL_n=logP\sum z_i+(1-\sum z_i)log(1-p)"


Thus "\dfrac{dL_n}{dp}=0"


"=\dfrac{\sum z_i}{p}-\dfrac{1}{1-p}\sum (n-\sum z_i)=0"


"=\dfrac{1}{P}\sum z_i"


"=\dfrac{1}{1-P}(n-\sum z_i)=n"


"\Rightarrow \hat{p}=\dfrac{1}{n}\sum z_i=\bar{z}"

This is the required MLE.